The period of a mass-on-spring oscillator is 3 seconds. Upon decreasing its mass by 500 g, the period will become 2 seconds.

a) Calculate the original mass!

b) Calculate the spring constant!


1
Expert's answer
2021-06-13T17:35:30-0400
"T=2\pi\sqrt{\frac{m}{k}}\\\frac{3}{2}=\sqrt{\frac{m}{m-0.5}}\\m=0.9\ kg"

"3=2\pi\sqrt{\frac{0.9}{k}}\\k=4\frac{N}{m}"